OpenStudy (anonymous):

Linear Inequalities: Sally intends to give 12 good friends toys. Her friends can either have a teddy bear or a toy car. A teddy bear cost $1.50 while a toy car cost $1.20. If she only has $16 to spend on the gifts, what is the max n.o of teddy bear her friend can have. Solve in linear inequalities.:D Let x be the max n.o of teddy bear This is what i have done so far 1.50x-1.2(12-x)<16 2.7x<30.4

OpenStudy (callisto):

Let x be the no of teddy bear. Since her friend can either have a teddy bear or a car. The number of car she buys = 12-x For the total price of buying cars, it is = 1.2 (12-x) For the total price of buying teddy bear, it is = 1.5x So the total money she spends on the gifts = total price of buying cars + total price of buying teddy bear = 1.2(12-x) +1.5x Note that she has only 16 dollars, so the amount she spends cannot exceed $16 So, the total amount spent should be less than or equal to 16, and you can write the inequality like this: total price of buying cars + total price of buying teddy bear ≦ 16 1.2(12-x) +1.5x ≦ 16 14.4 -1.2x +1.5x ≦ 16 Can you solve it from here? Note that once you find that it is a fraction ( or decimal) , say 5.23, that maximum no of teddy bear you can buy is the largest integer satisfying the inequality, that is 5 (5.23 is not the answer, it's just an example) Can you work it out now?

OpenStudy (anonymous):

b + c = 12 toys 1.5b + 1.2 c <= $16 -> 1.2b + 1.2c = 14.4 => b = 1.6 / .3 = 5.3 Max is 5 bears